OpenStudy (anonymous):

i need help with calculus. find the limit as x approaches 1 . x-1/|x-1|

OpenStudy (anonymous):

OKay Limit doesn't exist as we approach from left we get minus inside the modulus making it -(x-1) and when we approach from right we have + inside mod x+1 so limit is 1 and -1 ..hence limit doesn't exist

OpenStudy (anonymous):

okay this makes sense, but how could you show work? do you plug in 1?

OpenStudy (anonymous):

myininaya is here she is better at explaining

myininaya (myininaya):

satellite is suppose to make is pretty piecewise function but hes not coming :( so i guess i will explain without his pretty piecewise function

OpenStudy (anonymous):

hmm you just put number greater than 1 a little greater than 1 when you approach from right side and a little smaller when you approach from left side hmm you can check khan academy they have some pretty nice videos

myininaya (myininaya):

\[x-1>0=> |x-1|=x-1; x-1<0=> |x-1|=-(x-1)\]

myininaya (myininaya):

\[\lim_{x \rightarrow 1^+}f(x)=\lim_{x \rightarrow 1^+}\frac{x-1}{x-1}=1\] \[\lim_{x \rightarrow 1^{-}}f(x)=\lim_{x \rightarrow 1^-}\frac{x-1}{-(x-1)}=-1\]

OpenStudy (anonymous):

By definition of absolute value when you approach from the right you have\[y=\frac{x-1}{x-1}\]since the abs value is positive, but when you approach from the left you have\[y=\frac{x-1}{-(x-1)}\]

myininaya (myininaya):

but \[1 \neq -1 \]

myininaya (myininaya):

so actual limit does not exist

OpenStudy (anonymous):

thank you so much the illustrations made it better